OpenStudy (anonymous):

How much work has the battery done from the time the switch was closed until t = 6.51×10-3 s?

OpenStudy (anonymous):

Relevant information: The resistance in the circuit is R = 140 Ω, the inductance is L = 2.40 H, and the battery maintains a voltage of ℰ = 30.0 V. At time t=0 the switch is closed. So I know I have to integrate the power formula P=IV =I ℰ where I= \[ (ℰ /R)(1-e ^{-Rt/L}) \] which result in the integrand \[\int\limits_{0_{}}^{t_0}ℰ(ℰ /R)(1-e ^{-Rt/L}) dt\] then simplified I got \[(ℰ^{2}/R ^{2})(t _{0}R +Le^{-Rt/L} -L) \] just wondering where I might have gone wrong with this integrand

OpenStudy (michele_laino):

the equation of your circuit is: \[\large E = RI + L\frac{{dI}}{{dt}}\] whose solution, using your initial conditions, is: \[\large I\left( t \right) = \frac{E}{R}\left[ {1 - \exp \left( { - \frac{R}{L}t} \right)} \right]\] Now, if we can neglect the dissipation of energy due to Joule effect, then the power of your battery is: \[\large \int\limits_0^{{t_0}} {EI\left( t \right)dt} = \frac{{{E^2}}}{{{R^2}}}\left[ {R{t_0} + L\exp \left( { - \frac{R}{L}{t_0}} \right) - L} \right]\] So, your answer is right!

OpenStudy (irishboy123):

.